### v1

# QCD Phase Diagram and the Finite Volume Fireball: A Model Study

Abstract
Experimental investigations of the phase diagram of strongly interacting matter involve collisions of heavy ions at ultrarelativistic velocities. The medium created in such a collision is often of dimensions a few fermi, in particular in the Beam Energy Scan experiments. An understanding of the effect of the finite volume and the boundary is important for connecting the experimental results to the phase diagram.
Using the Nambu Jona-Lasinio model, an effective theory for the chiral transition of quantum chromodynamics (QCD), we have studied the effect of the finite volume of the fireball on the transition line at finite temperature and density using the MIT boundary condition to mimic the condition that the system is deconfined inside. The shift of the transition temperature for finite volume and finite volume effect on number density and its susceptibilities are studied. The volume effects should be considered when looking for signatures of the phase diagram in experiments.
Finite volume, QCD phase diagram, NJL model, MIT boundary condition.
I Introduction
The bulk property of the strongly interacting matter has been a subject of detailed experimental investigation in recent times [1, 2, 3, 4]. Strongly interacting matter is expected to manifest itself in a deconfined, chiral symmetry-restored phase, called the quark-gluon plasma (QGP), at very high temperatures [5, 6]. Experimentally, this phase has been explored in the study of the medium created in the collision of heavy ions at ultrarelativistic speeds, in particular in the Relativistic Heavy Ion Collider (RHIC) in the USA [3] and the Large Hadron Collider (LHC) in CERN, Geneva [1]. The phase structure at high number density is being explored in the Beam Energy Scan (BES) experiments in RHIC [4] and will be further explored in future experiments in NICA, JINR [7]. The knowledge about the phases at high density is complemented by the data from neutron stars and their mergers [8]. On the theory side, the properties of the deconfined state at high temperatures are well-understood from large-scale numerical studies of quantum chromodynamics (QCD), the theory of strongly interacting matter, using lattice discretization [5]. Lattice studies are more difficult at nonzero baryon number chemical potential ($\mu_{\scriptscriptstyle B}$); however, the equation of state and baryon number susceptibilities have been calculated at large temperatures and moderate values of $\mu_{\scriptscriptstyle B}$. At higher values of $\mu_{\scriptscriptstyle B}$, or lower temperatures, one has to rely on QCD-based models [9, 10].
The theoretical studies, in particular the lattice-discretized QCD (LQCD) studies, are for static equilibrated infinite volume systems. On the other hand, the medium created in RHIC and LHC experiments is rapidly evolving, and the system size is not too large. In particular, in the BES studies the system dimension can be a few fermi, not too large compared to the typical correlation length in strong interaction physics. The current experimental efforts are to study QGP in smaller systems. All this makes an understanding of the volume effect important.
In this paper, the aim is to get an idea of the effect of the finite volume on the phase diagram for a static system, with the effects of the dynamic nature left for a future study. We study the phase diagram using the Nambu Jona-Lasinio (NJL) model [9], which is an effective model expected to capture the physics related to the chiral symmetry, in particular, the spontaneous chiral symmetry breaking. Since the finite temperature crossover is associated with the melting of the chiral condensate, the NJL model is expected to capture the basic features of the crossover.
In a study of the finite volume effects it is essential that realistic boundary conditions are put in for the system. The actual boundary of the medium created in the heavy ion collision experiments is of complicated geometry. We take the simplified MIT boundary condition [11] (see [12] for a pedagogic discussion). While simple, the MIT boundary condition captures the essential physics that inside the boundary we have a deconfined medium.
The effect of the finite volume on the chiral transition has been studied before [13, 14, 15, 16, 17, 18, 19, 20]. These studies have used the anti-periodic (APBC) and periodic (PBC) boundary conditions, or have put in a lower cutoff in momentum. However, these boundary conditions do not mimic a finite-volume fireball; they are boundary conditions used in theoretical studies to mimic an infinite volume system within a finite-volume setup, e.g., in numerical lattice studies. In the Appendix B we compare the MIT boundary condition results with those obtained with other geometries. The MIT boundary condition has been used before, for studying the chiral crossover at $\mu_{\scriptscriptstyle B}=0$ on a sphere [21], and for a rotating fermionic system in a cylinder [22, 23]. Here we use it to study the phase diagram in the $\left(T,\,\mu_{\scriptscriptstyle B}\right)$ plane and the baryon number susceptibilities.
The plan of the paper is as follows. In Section II we outline the basic calculations, discuss the boundary conditions, and specify the parameters used. The change of the phase diagram with changing volume is discussed in Section III. The number density and the susceptibilities in the critical region are discussed in Section IV. We summarize our results in Section V. The relevant equations for the MIT boundary condition and other details are compiled in Appendix A. Appendix B has a discussion of other boundary conditions.
II Formalism and the boundary condition
The NJL model is an effective model of the spontaneous chiral symmetry breaking of QCD [24, 9]. The degrees of freedom are only quarks; the gluonic interactions are included in four-fermion interaction terms, chosen such that the Lagrangian is chirally symmetric.
We will work with the two-flavor, isospin symmetric theory. The NJL Lagrangian for two degenerate flavors of light quarks is,
$\mathcal{L}_{\scriptscriptstyle\rm NJL}={\bar{\psi}}(i\gamma_{\mu}\partial^{\mu}-{\hat{m}})\psi+G\left[({\bar{\psi}}\psi)^{2}+({\bar{\psi}}i\gamma_{5}{\bar{\tau}}\psi)^{2}\right],$ (1)
where $\psi=\begin{pmatrix}\psi_{u}\\ \psi_{d}\end{pmatrix}$ is the quark field, ${\hat{m}}=\text{diag}(m,m)$ is the degenerate light quark mass matrix, and $G$ is the coupling strength of the effective four-fermion interaction between quarks. In the massless limit ($m=0$), $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ has a global $SU(2)_{L}\times SU(2)_{R}\times U(1)_{V}$ symmetry. This symmetry is spontaneously broken to $SU(2)_{V}\times U(1)_{V}$ as the chiral condensate $\langle\bar{\psi}\psi\rangle$ takes a nonzero value. In the presence of the mass term, the symmetry is also explicitly broken, leading to a preferred vacuum in the broken phase.
We will be studying the theory in the mean-field (MF) approximation. Writing the chiral condensate to be:
$\langle\bar{\psi}\psi\rangle\;=\;-\frac{\sigma}{2G},$ (2)
$\mathcal{L}_{\scriptscriptstyle\rm NJL}$ can be rewritten in the mean-field approximation as,
$\mathcal{L}^{\scriptscriptstyle\rm MF}_{\scriptscriptstyle\rm NJL}\;=\;\bar{\psi}(i\gamma_{\mu}\partial^{\mu}\,-\,M)\psi\,-\,\frac{\sigma^{2}}{2G},\qquad M\,=\,m\,+\,\sigma\,.$ (3)
Then the effective potential is given by,
$V_{\rm eff}(\sigma)\;=\;-\frac{1}{V}\,\log Z\;=\;\frac{\sigma^{2}}{2G}\,-\,\int\mathbb{D}p\,\log\,\det\left(\gamma\cdot p\,-\,M\right)\,,$ (4)
and $\sigma$ can be obtained from the minimum of $V_{\rm eff}(\sigma):\ \frac{\textstyle\partial V_{\rm eff}}{\textstyle\partial\sigma}\,=\,0.$ Here $\mathbb{D}p$ refers to the regularized momentum integral.
The four-fermion term in $\mathcal{L}_{\scriptscriptstyle\rm NJL}$ makes the theory non-renormalizable. Hence, a suitable regularization method has to be specified, and the results will depend on the regularization parameter. We have used Schwinger’s proper time regularization [24, 25], with the cutoff $\tau_{\rm\scriptscriptstyle UV}=1/\Lambda_{\rm\scriptscriptstyle UV}^{2}$. Therefore, our model has three parameters: $m,G$ and the UV momentum cutoff $\Lambda_{\rm\scriptscriptstyle UV}$. These parameters can be chosen by specifying the pion mass, the pion decay constant and either the chiral condensate or the chiral transition temperature ($T_{c}$). Some suitable parameter sets, and how they affect the chiral crossover temperature, have been discussed in Ref. [25]. The phase diagram can be qualitatively different for the different parameter sets. In particular, for some choices of parameters one gets a critical point and a first-order transition at large quark number chemical potential $\mu_{q}=\dfrac{\mu_{\scriptscriptstyle B}}{3}$, while for some other choices, the transition is a crossover in the whole $\left(T,\,\mu_{q}\right)$ plane. We have chosen the following parameter set:
$m$ (MeV) $G$ (GeV-2) $\Lambda_{\rm\scriptscriptstyle UV}$ (MeV)
15 17.2 645
This parameter set was already explored in Ref. [25]. It has a first-order transition line at large $\mu_{q}$, ending at a critical point at $\mu_{q,c}\approx$ 350 MeV. At $\mu_{q}=0$, it shows a chiral transition at $T_{\chi}\approx$ 186 MeV. Both the $T_{\chi}$ and the $\mu_{q,c}$ are larger than the expected values for QCD. However, here our aim is not to tune the absolute value of these parameters but to look at their variation with the system size.
We will be studying the system at finite temperature $T$ and quark number chemical potential $\mu_{q}$. $\mu_{q}$ is introduced by adding $\mu_{q}\psi^{\dagger}\psi$ to Eq. (1). To get the thermodynamic potential, the integral over $p_{0}$ in Eq. (4) has to be turned into a fermionic Matsubara sum. For finite volume, the integral over spatial momenta, $\dfrac{d^{3}p}{(2\pi)^{3}}$, will be replaced by a sum over the momentum modes allowed by the boundary condition. For a study of the effect of finite volume, imposing a proper boundary condition is important. As we discussed in Section I, the exact boundary condition for the medium created in RHIC or LHC is complicated; since it is time-dependent and also changes from event to event. Since we are only interested in getting an idea about how different the finite volume system can be from an infinite volume system, we will take a simple spherical geometry. The important physics that we want to capture is that the deconfined QGP is only contained within this geometry. For this, we impose the MIT boundary condition, which implies that the normal component of the quark current becomes zero at the boundary. More precisely, we impose the boundary condition [11, 12],
$(-i{\hat{r}}\cdot{\gamma})\psi(t,r,\theta,\phi)|_{r=R}=\psi(t,r,\theta,\phi)|_{r=R},$ (5)
where $\hat{r}$ is the unit vector normal to the spherical surface and $\gamma_{\mu}$ are the Dirac matrices. Eq. (5) ensures the vanishing of the normal component of the fermionic current and scalar densities of quarks at the boundary of the spherical surface:
$J_{n}\equiv n_{\mu}\bar{\psi}\gamma^{\mu}\psi=0,\qquad\text{and}\qquad\bar{\psi}\psi=0\ \ \ \text{at}\ \ \ r=R,$ (6)
where $\hat{n}$ is the unit normal four-vector to the boundary. The allowed momentum modes for Eq. (5) are well-known; the relevant equations are given in Appendix A. Then the expression for $\Omega$ becomes:
$\Omega=\frac{(M-m)^{2}}{4\,G}-\frac{N_{c}\,N_{f}}{V}\sum_{j}\sum_{p_{j_{\kappa}}}(2j+1)\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu_{q})}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu_{q})}{T}}\right)\right],$ (7)
[…]
Effects of finite volume on the phase diagram have been explored in the literature in the past. Sometimes the antiperiodic or the periodic boundary conditions have been used for this purpose [15, 16, 26, 19]. Note that these boundary conditions do not capture the finite volume fireball in any way: these are boundary conditions usually imposed in theoretical studies (e.g. numerical studies on the lattice) when the aim is to study the infinite volume system. Besides being easy to implement, the advantage of these boundary conditions is that they show very small deviations from the infinite volume system, and are therefore ideal for studies aiming for the thermodynamic limit. Our purpose here is to estimate the size of the deviation from the infinite volume results due to the finite volume fireball. We do not think (anti)periodic boundary conditions are suitable for this purpose. While the boundary condition Eq. (5), Eq. (6) also do not capture the whole complexity of the finite fireball, we believe it captures the most essential aspects of it and gives a more realistic estimation of the size of the finite volume effects. In Appendix B we compare our results with results with an antiperiodic geometry for a similar system size. We indeed find that the volume effects are much smaller in the antiperiodic geometries.
The finite temperature phase transition is chiral symmetry restoring in nature, and we will track the transition using $\langle\bar{\psi}\psi\rangle$, or equivalently, $M$. In the absence of the symmetry-breaking term $m$, in the full quantum theory, the chiral symmetry will be restored in finite volume at all temperatures, due to tunnelling between the classical vacua of $\mathcal{L}^{\scriptscriptstyle\rm MF}_{\scriptscriptstyle\rm NJL}$. In Monte Carlo simulations, this will show an instability in $\langle\bar{\psi}\psi\rangle$. A mean-field treatment will miss the physics of tunnelling, and indicate symmetry breaking at low temperatures. This problem is not there in the presence of the mass term corresponding to physical light quarks, where lattice simulations show a stable result for $\langle\bar{\psi}\psi\rangle$ at quite small volumes down to $m_{\pi}L\sim 1$.
III Chiral transition at finite volume
We explore the chiral transition line for Eq. (3), and its shift with volume, for the MIT boundary condition. At $\mu_{q}=0$ the chiral symmetry restoration is a crossover. In Fig. 1, we show the constituent mass, $M$ (Eq. (10)), which is just a constant shift of $-\langle\bar{\psi}\psi\rangle$ and has the same information as $\langle\bar{\psi}\psi\rangle$. We see a rapid but smooth decrease in $M$ around $T\approx$ 190 MeV, indicating a crossover. At $T=0$, decreasing the volume to $R=5$ fm decreases the constituent mass by $<5\%$ from the infinite volume value. At high temperatures $>200$ MeV the chiral symmetry is restored and the chiral condensate becomes small. Therefore $M$ becomes small, approaching $m$. In small volume systems, the crossover region becomes somewhat smoother, but no qualitative change from the infinite volume case is seen in systems as small as a 5 fm sphere. This can be further seen in the right panel of Fig. 1, where the derivative $dM/dT$ is shown. To get this plot, we used a cubic spline interpolation of the $M$ vs $T$ data of the left plot. The plot shows a systematic but moderate reduction in the peak height and broadening of the peak, as the volume is reduced. The peak position is an estimate of the crossover temperature $T_{c}$, which changes very slowly with the radius of the sphere, $R$ (a $7$ MeV shift from infinite volume to $R=5$ fm).
Figure 1: (Left) Constituent quark mass $M$ as a function of temperature ($T$) at $\mu$ = 0. (Right) The derivative $dM/dT$. The legend shows the peak position of the $dM/dT$ curves, which can be taken as an estimate of $T_{c}$.
Instead of $dM/dT$, it is more common to look at the chiral susceptibility $\chi_{m}=-\frac{\textstyle\partial\langle\bar{\psi}\psi\rangle}{\textstyle\partial m}$. The chiral susceptibility shows a peak at the chiral symmetry restoration point. The peak position is close to the peak of $dM/dT$, though not exactly identical at the crossover region, and the volume dependence is similar to that seen in Fig. 1. Here we will use the peak of $dM/dT$ as our estimate for $T_{c}$. Of course, in the case of a true phase transition (as opposed to a crossover), all these observables will show the same transition temperature.
Now we introduce a nonzero quark chemical potential $\mu_{q}$. For small values of $\mu_{q}$, the qualitative behavior is very similar to what is seen for $\mu=0$. In Fig. 2, we show $M$ as a function of $T$ at different volumes for three values of $\mu_{q}$. The temperature and volume dependence of $M$ at $\mu_{q}$=100 MeV is very similar to that at $\mu$=0, as a comparison of the left panel of Fig. 2 with Fig. 1 shows. As $\mu_{q}$ increases, the crossover shifts to smaller values of temperature, as expected. Also, we find a stiffening of the transition at larger volumes: the smoothing of the transition at small volumes is enhanced. At $\mu_{q}$=350 MeV, shown in the right panel of Fig. 2, we find a discontinuity, indicating a first-order transition, at infinite volume. The first-order nature of the transition can be further ascertained by looking at $\Omega$, Fig. 3. At the transition region, a two-minimum structure is seen, which is the hallmark of first-order transition, and indicates phase coexistence. At this $\mu_{q}$ the transitions are smooth in the finite volume system, even at $R$=20 fm, though at such a large volume it is quite a sharp crossover. We also find here a much more enhanced smoothing of the transition with a reduction of volume: while the curve at $R$ = 20 fm indicates a sharp crossover, the one at $R$=5 fm indicates a smooth behavior. Also, the crossover temperature changes much more rapidly here with volume: the transition temperature changes from 47.3 MeV in the infinite volume system to 72.2 MeV in the 5 fm sphere.
Refer to caption Refer to caption Refer to caption
Figure 2: The dependence of the $M$ vs $T$ curve with volume. Results at $\mu_{q}$= 100 MeV (left), 250 MeV (middle) and 350 MeV (right) are shown. A discontinuous transition is seen in the infinite volume system at $\mu_{q}$=350 MeV.
Figure 3: The thermodynamic potential, $\Omega/\Lambda_{\rm\scriptscriptstyle UV}^{4}$, in the transition region at $\mu_{q}$=350 MeV for an infinite volume system. A small (irrelevant) additive constant $\Omega=\mp 0.0001\,\Lambda_{\rm\scriptscriptstyle UV}^{4}$ has been added to the potentials at T=44 and 50 MeV, respectively, for better viewing. The figure indicates a first-order transition at $T\approx$ 47.3 MeV.
Before proceeding with the discussion of the investigation of the transition line at higher $\mu_{q}$ at finite volume, let us clarify our terminology. In our analysis, the term ”first-order transition” corresponds to a two-well structure in the free energy, with well-separated local minima at small and large values of $M$ corresponding to the symmetry restored and symmetry broken phases, respectively. Such a structure will indicate phase coexistence, and possible supercooling/superheating in a dynamical system. This is what we have looked for, and have referred to as the signature of a first-order transition.
In the infinite volume case, tunnelling between the two minima is prohibited. However, in a finite volume, the free energy barrier between the minima is finite, and there is a nonzero probability for tunnelling. This will lead to a unique minimum at every $\left(T,\,\mu_{q}\right)$ value for a static system, and strictly speaking, one can only have crossovers in the finite volume system. However, in a fireball which rapidly cools through the transition region, tunnelling will be extremely suppressed. In cases where the classical free energy shows a two-minima structure and a discreet jump at the transition, we will expect interesting phenomena, related to the features of a first-order transition, at the transition point. Therefore, the two-well structure of the mean-field potential, $\Omega_{\scriptscriptstyle MF}$, is of actual physical interest, and we will loosely use terms like first-order line, phase coexistence etc. to refer to such a structure even in a finite volume system.
Refer to caption Refer to caption Refer to caption
Figure 4: The mean-field thermodynamic potential in the transition region at $T$=20 MeV for infinite volume (left) and for spheres of radius 20 fm (middle) and 10 fm (right), showing a two-well structure down to a sphere of $R$=10 fm.
We illustrate this behavior in Fig. 4, where we show $\Omega_{\scriptscriptstyle MF}$ at different volumes at $T$=20 MeV. We find a two-well structure in $\Omega_{\scriptscriptstyle MF}$ even at $R$=10 fm. Note also the range of $\mu_{q}$ over which this structure is seen (we loosely refer below to this region as the transition region). A shift of the transition to higher $\mu_{q}$ is seen with the reduction in volume; at the same time, a weakening of the double-well structure (and a shrinking of the transition region) is seen. At this temperature, we do not find a discontinuous transition in the 5 fm sphere, but as we go further down to $T$=0 we find a discontinuous transition in all our systems.
Our finding of the transition line in the $(T,\mu_{q})$ plane is summarized in Fig. 5. In the inset, the first-order parts of the line are shown. Here the dotted lines indicate the estimate of the boundary of the two-phase region.
Figure 5: The shift of the chiral crossover line in the $(T,\mu_{q})$ plane with system size. In the inset the large $\mu_{q}$ part is enlarged, highlighting the first-order region.
IV Number density and quark number susceptibility
The chiral condensate $\langle\bar{\psi}\psi\rangle$ and, equivalently, the constituent mass $M$ are good order parameters for tracking the chiral transition. They are, however, not directly connected to an experimental observable. The baryon number density ($n_{\scriptscriptstyle B}$) and susceptibilities of the baryon number operator are, on the other hand, related to experimental observables. In particular, they are of great interest in the large $\mu_{\scriptscriptstyle B}$ regime.
[…]
$n_{q}=\frac{\partial P}{\partial\mu_{q}}=2N_{c}N_{f}\int\frac{d^{3}p}{(2\pi)^{3}}\left[n_{\scriptscriptstyle F}(E-\mu_{q};T)-n_{\scriptscriptstyle F}(E+\mu_{q};T)\right]\,,$ (8)
where $n_{\scriptscriptstyle F}$ is the Fermi distribution, $n_{\scriptscriptstyle F}(x;T)=\left(e^{\tfrac{x}{T}}+1\right)^{-1}$.
We start with an investigation of the effect of finite volume on $n_{q}$ in the crossover region. In the left panel of Fig. 6, we show the temperature dependence of $n_{q}$ at $\mu_{q}$=100 MeV. At small values of $\mu_{q}\lesssim 100$ MeV, the volume dependence of number density is small except in the crossover region (150 MeV $\leq\ T\ \leq$ 200 MeV). $n_{q}$ is small at small temperatures, and then rises and reaches its asymptotic behavior $\sim\mu_{q}T^{2}$ at high temperatures. In the crossover zone, we see some volume dependence: the rise of $n_{q}$ is milder at smaller volumes. This behavior gets further amplified as we go to higher values of $\mu_{q}$, where the transition becomes sharper. In the middle panel of Fig. 6 we show the results for $\mu_{q}$=250 MeV, still in the crossover region. Since the crossover is now sharper, $n_{q}$ increases much more sharply across the transition line at larger volume systems. Also above the transition the temperature is still not sufficiently high to show the asymptotic $\sim\mu T^{2}$ behavior. As we go to even higher values of $\mu_{q}$, we reach the first-order region, where $n_{q}$ shows a discontinuous jump across the transition. This is seen in the infinite volume case of the right panel of Fig. 6 ($\mu_{q}$=350 MeV). Also here the behavior of $n_{q}$ above the transition is quite different, as the $\mu_{q}^{4}$ term starts dominating.
Refer to caption Refer to caption Refer to caption
Figure 6: The quark number density ($n_{q}$) as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu_{q}$ of 100 MeV (left), 250 MeV (middle) and 350 MeV (right).
The quark number density, and the nonlinear quark number susceptibilities which are higher-order derivatives of the free energy, are very interesting observables, especially for the critical point search. The nonlinear susceptibilities are defined as derivatives of the pressure ($P$) with respect to the quark chemical potential. Similar susceptibilities of other conserved quantities like charge or strangeness can also be constructed, but here we will concentrate on the quark number susceptibilities or equivalently, baryon number susceptibilities. The $n$th order susceptibility $\chi_{n}$ is defined as,
$\chi_{n}=\frac{\partial^{n}\,P}{\partial\,\mu_{q}^{n}}=\frac{\partial^{n-1}\,n_{q}}{\partial\,\mu_{q}^{n-1}}\,.$ (9)
The second-order susceptibility $\chi_{2}$ is a measure of the fluctuation of the baryon number. It shows a peak in the transition region. Near the critical region, the peak height becomes large, diverging at the critical point in the thermodynamic limit. In a finite volume system, the peak height scales with volume. In Fig. 7, we have shown $\chi_{2}$ at various values of $\mu_{q}$. Note the rapid increase in peak height and sharpened peak structure at the higher values of $\mu_{q}$.
Refer to caption Refer to caption Refer to caption
Figure 7: The quark number susceptibility $\chi_{2}$ as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu_{q}$ of 100 MeV (left), 250 MeV (middle) and 350 MeV (right).
It has been suggested [27, 28] that the higher-order susceptibilities are a better probe for the critical region. In particular, $\chi_{3}$ [29] and $\chi_{4}$ [30, 31] show very interesting structures in the critical region, and these structures become smoother rapidly with decreasing volume. $\chi_{3}$ has a peak to the lower $\mu_{q}$ side of the transition line, then changes sign and shows a dip in the higher $\mu_{q}$ side. These peaks and dips become sharper as one approaches the critical point.
V Summary of results and conclusion
The QCD phase diagram has been the subject of intense theoretical and experimental investigation for a few decades. The phases at small values of $\mu_{\scriptscriptstyle B}$ and the crossover line are well-understood from non-perturbative numerical studies of QCD using lattice discretization. The phase diagram at large values of $\mu_{\scriptscriptstyle B}$ has to be studied using effective models of QCD. In particular, the Nambu Jona-Lasinio model in its different variations has been widely used to study the phase diagram in the $\left(T,\,\mu_{\scriptscriptstyle B}\right)$ plane. On the experimental side, the phase diagram is studied by creating a strongly interacting medium at high temperatures and chemical potential in ultrarelativistic heavy ion collisions, in particular, in the BES series of experiments at RHIC, and other planned experiments like the NICA at JINR. To infer the phase diagram from the experiments, it is necessary to compare observables, in particular the number density and its susceptibilities, with theoretical calculations.
The fireball created in the experiments has a finite volume. It is rapidly evolving in time, and its dimensions are a few fermi before freezeout. On the other hand, the theoretical studies are usually for infinite volume, static systems. To infer information about the phase diagram by combining the two, it is therefore important to estimate the effect of the finite size and the dynamic nature of the fireball on the observables studied. In this paper, we have studied one aspect of it, the effect of the finite size of the system. We have studied the 2-flavor NJL model on a sphere, with MIT boundary conditions to mimic the fireball property of having a deconfined medium within the finite volume. While this is an idealization of the complicated geometry of the fireball created in a heavy ion collision, the MIT boundary condition captures the essential feature of the finite volume fireball that the quarks are confined within a finite system. Therefore one can reasonably expect to learn about the effect of the finite size using this setup.
The result of our investigation can be summarized by Fig. 5, which shows the shift of the transition line with the system volume. For the parameter sets used in our study, the NJL model has a first-order transition line at large $\mu_{q}$, ending in a critical point at $\left(T,\,\mu_{q}\right)\approx$ (60, 340) MeV. The actual location of the transition line as well as the endpoint is sensitive to the parameters chosen; the interest of this paper is on the sensitivity of the line on the system volume. Here the transition line has been tracked by the constituent quark mass $M$, or equivalently, $\langle\bar{\psi}\psi\rangle$. As the figure shows, in the crossover region, the shift of the transition line is mild: even for a system with $R$=5 fm, the shift is only a few MeV. The change of the number density with volume is shown in Fig. 6. As expected, the change of the number density across the crossover line becomes smoother as the volume is reduced.
The effect of the finite volume on the transition line is more interesting in the first-order region on the high $\mu_{q}$ side. The first-order region (indicated by the two-minima structure of the mean-field free energy; see Section 3) ends at a lower value of $T$ as the volume is reduced. The number density also shows very similar behavior. The volume dependence is particularly strong in the quark number susceptibilities of the baryon number in the critical region, as seen in Fig. 7. One would expect even larger effects for the higher-order susceptibilities. These results will be relevant for the experimental searches for the critical region, where the critical region is looked for in a small fireball.
VI Acknowledgement
A.S. thank Prof. M. Chernodub for the helpful discussion. The computing resources of the Department of Theoretical Physics, TIFR, were used for this work. We would like to thank Ajay Salve and Kapil Ghadiali for their technical assistance. S.D. acknowledges the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002.
Appendix A Calculational details
In this section, we give some formulae used in the calculations for obtaining results of Section III and Section IV, and some relevant details.
The chiral condensate causes the quarks to have an effective mass $M$, and at the mean-field level, the theory can be treated like a Gaussian theory with quarks of mass $M$. $M$ is calculated self-consistently using the following equations,
$M=m-2\,G\,\langle\bar{\psi}\psi\rangle,\qquad\qquad\langle\bar{\psi}\psi\rangle=-\int\mathbb{D}p\,\text{Tr}[S(p,M)],$ (10)
where $\mathbb{D}p$ is the regularized four-momentum integral and $S(p,M)$ is the free quark propagator matrix.
We work with the proper time regularization, in which the momentum regulator is imposed as,
[…]
and we get the chiral condensate,
[…]
We can easily isolate the terms quadratic and logarithmic in $\Lambda_{\rm\scriptscriptstyle UV}$:
$\langle\bar{\psi}\psi\rangle\ =\ -\frac{N_{f}\,N_{c}\,M^{3}}{4\pi^{2}}\left(\frac{\Lambda_{\rm\scriptscriptstyle UV}^{2}}{M^{2}}\,e^{-M^{2}/\Lambda_{\rm\scriptscriptstyle UV}^{2}}\;-\;\Gamma\left(0,\,\frac{M^{2}}{\Lambda_{\rm\scriptscriptstyle UV}^{2}}\right)\right),\qquad\Gamma(0,x)\,=\,-\gamma\,-\,\log x\,-\,\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\,\cdot\,k!}\,,$ (13)
where $\Gamma(0,x)$ is the incomplete Gamma function, and $\gamma\approx 0.577$ is the Euler’s constant [32].
$\mathcal{L}_{\scriptscriptstyle\rm NJL}$ is a non-renormalizable theory, and the results depend on $\Lambda_{\rm\scriptscriptstyle UV}$. A comparison of various regularization schemes can be found in Ref. [25]. In particular, the result for $\langle\bar{\psi}\psi\rangle$ in dimensional regularization with $\overline{MS}$ scheme is,
$\langle\bar{\psi}\psi\rangle_{\scriptscriptstyle\overline{MS}}=-\frac{N_{f}\,N_{c}\,M^{3}}{4\pi^{2}}\left(-1\,-\,\log\frac{\mu^{2}}{M^{2}}\right),$ (14)
where $\mu$ is the $\overline{MS}$ scale, and we have done the standard $\overline{MS}$ subtraction $\dfrac{1}{\epsilon}-\gamma+\log(4\pi)$ from the right hand side of Eq. (14). If we were to do the standard renormalization prescription of taking $\Lambda_{\rm\scriptscriptstyle UV}\to\infty$ in Eq. (13), then matching to Eq. (14) would require a subtraction of
$-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\Lambda_{\rm\scriptscriptstyle UV}^{2}\,+\,M^{2}\left(\log\frac{\mu^{2}}{\Lambda_{\rm\scriptscriptstyle UV}^{2}}\,+\gamma\right)\right]$
at this order. Of course, here $\Lambda_{\rm\scriptscriptstyle UV}$ is kept finite, and the result depends on the details of the regularization.
At finite temperature, the integral over $p_{0}$ gets replaced by a sum over the Matsubara modes $(2n+1)\pi T$, leading to the mean-field result for the thermodynamic potential:
$\Omega=\frac{(M-m)^{2}}{4\,G}-2\,N_{c}\,N_{f}\int\dfrac{d^{3}p}{(2\pi)^{3}}\left[\left\{-\frac{\Lambda\,e^{-\left(\frac{E}{\Lambda}\right)^{2}}}{\sqrt{\pi}}+E\,\text{Erfc}\left(\frac{E}{\Lambda}\right)\right\}+T\,\text{log}\left(1+e^{-\frac{(E-\mu_{q})}{T}}\right)+T\,\text{log}\left(1+e^{-\frac{(E+\mu_{q})}{T}}\right)\right].$ (15)
The chiral condensate becomes:
$\langle\bar{\psi}\psi\rangle(T,\mu_{q})\,=\,-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\int_{\tau_{\rm\scriptscriptstyle UV}}^{\infty}\frac{d\tau}{\tau^{2}}\,e^{-M^{2}\tau}\;-\;\frac{1}{E_{p}}\left(n_{\scriptscriptstyle F}(E_{p}-\mu;T)\,+\,n_{\scriptscriptstyle F}(E_{p}+\mu;T)\right)\right],$ (16)
where $E_{p}=\sqrt{p^{2}+M^{2}}$. Note that we have only regularized the vacuum term, as the terms with the Fermi distributions are already regularized by the temperature scale.
The momentum modes for the MIT boundary condition on a sphere can be obtained by solving the Dirac equation for fermions of mass $M$ and imposing the MIT boundary condition Eq. (5) on the eigenfunction. The allowed momentum modes satisfy the condition [12],
$j_{l_{\kappa}}(pR)=-\,\text{sign}(\kappa)\left(\frac{p}{E+M}\right)j_{l_{\bar{\kappa}}}(pR).$ (17)
Here, $j_{l}$ is $l^{th}$ order spherical Bessel function, and
$l_{\kappa}=\begin{cases}-\kappa-1&\quad\text{for}\quad\kappa<0\\ \kappa&\quad\text{for}\quad\kappa>0\end{cases}$ (18)
$l_{\bar{\kappa}}=\begin{cases}-\kappa&\qquad\text{for}\quad\kappa<0\\ \kappa-1&\qquad\text{for}\quad\kappa>0.\end{cases}$ (19)
The energy eigenvalue for such a system is,
$E\equiv E_{\kappa,i}=\sqrt{q_{\kappa,i}^{2}+M^{2}},$ (20)
where $q_{\kappa,i}\equiv pR$ is obtained by solving Eq. (17). Here, $i$ is the radial excitation quantum number ($i=1,2,3,...$) and $\kappa=\pm(j+1/2)$ is for the total angular momentum, $j=1/2,3/2,...$. Replacing the three-momentum integral in Eq. (15) by a sum over these modes leads to Eq. (7).
Appendix B Comparison with other boundary conditions
The QCD phase diagram at finite volume has been studied in the literature, using the NJL model [33, 21] and other effective models like the PNJL [13, 14], the Polyakov quark-meson model [17, 34], etc., as well as using the Dyson-Schwinger equation [26] and the functional renormalization group [15]. As we have mentioned in Section II, most of the studies have used the periodic (PBC) or antiperiodic (APBC) boundary conditions. For a cubic box of volume $L^{3}$, this corresponds to momentum sum with spatial momenta $\dfrac{2n\pi}{L}$ and $\dfrac{(2n+1)\pi}{L}$, respectively. Also in some works [13, 14, 17], an infrared cutoff $\dfrac{\pi}{L}$ has been used in the momentum integral to mimic a system of dimensions $L$, but the geometry has not been specified. As we have mentioned in Section II, we do not think such boundary conditions mimic the finite size of the fireball. While a gross simplification, the MIT boundary condition used here captures the essential physics of the finite-size fireball.
To see how much the results with the MIT boundary condition differ from the other boundary conditions, we have compared the size of the finite volume effects with the different boundary conditions, by looking at the constituent mass $M$. In Fig. 8, we have compared the results for the temperature dependence of $M$ for APBC and the IR cutoff for a box of dimension $L$=5 fm, with those of MIT boundary condition with $R$=5 fm.
Figure 8: Dressed quark mass ($M$) as a function of temperature ($T$) for different boundary conditions.
As anticipated in Section II, the finite volume effect is significantly larger with the MIT boundary than the other boundary conditions. The APBC and the IR momentum cutoff show very little deviation from the infinite volume result already at $L$=5 fm. The PBC is expected to show even less finite volume effect than the APBC, as has been seen in lattice studies. In the chiral symmetry broken phase, the finite volume effect with the MIT boundary condition on a sphere is clearly more than that with APBC and IR cutoff, despite the volume of the sphere being about four times larger than those with a cubic geometry. In the high-temperature symmetry restored phase the boundary effect is substantially smaller for all the boundary conditions. Fig. 8 indicates, in our view, that the studies with the APBC and similar boundary conditions significantly underestimate the finite volume effect.
While the APBC and PBC are not suitable for studying the finite volume fireball, they are excellent for numerical lattice studies, where one tries to get the infinite volume results from simulations on finite boxes. These boundary conditions are easy to implement on lattice, and show very little volume effect, making it easy to extrapolate to infinite volume from relatively small boxes. To check whether the mean-field NJL model qualitatively captures the volume dependence of the chiral condensate, we compared the results from lattice for $\langle\bar{\psi}\psi\rangle$ on small boxes with the NJL MF results. $\langle\bar{\psi}\psi\rangle$ is regularization dependent; in principle, one can convert them using results like those in Appendix A. Instead, we for simplicity compared the renormalization group invariant quantity:
$y=m\left[\langle\bar{\psi}\psi\rangle(L)-\langle\bar{\psi}\psi\rangle(L\rightarrow\infty)\right].$ (21)
The comparison of such a study is shown in Fig. 9. This is a qualitative comparison where the lattice pion is somewhat heavier ($\approx$ 200 MeV). Also for the lattice, the $L\to\infty$ is replaced by the largest lattice we used. So a precise agreement is not expected. But as Fig. 9 shows, the deviation from the full theory at a very small volume is qualitatively captured in the NJL MF result. This gives us hope that the results we obtained for the MIT boundary condition, despite the many caveats, qualitatively capture the trend of the finite volume effect of a static fireball.
Figure 9: Comparison of $y$, Eq. (21), between lattice and MF NJL with APBC.